OpenStudy (anonymous):

write an equation using two variables to show the relationship between the first number f and the second number s.

OpenStudy (anonymous):

pleas help me

jhonyy9 (jhonyy9):

like f=2s

jhonyy9 (jhonyy9):

than s=f/2

jhonyy9 (jhonyy9):

you have thought it in this way ?

jhonyy9 (jhonyy9):

or the same f=s+2 so than s=f-2

jhonyy9 (jhonyy9):

if those are right please push the good answer button

jhonyy9 (jhonyy9):

thank you

OpenStudy (anonymous):

>_< thanks